OpenStudy (emogirlatemoooocow):

Which points are 17 units from the point located at 46 on a number line? A. 29 and 63 B. –29 and 29 C. –29 and 63 D. –63 and 63

OpenStudy (anonymous):

A. {46+17=63,46−17=29}

OpenStudy (gorv):

there will be two point one on left and one on right

OpenStudy (gorv):

46+17 =63 and 46-17=29

OpenStudy (anonymous):

You got it?